OpenStudy (anonymous):

.06-.01(x+1)=.02(2-x) i got 294/100 ?

OpenStudy (lalaly):

.06-0.01x-0.01= 0.04-0.02x .05-.01x=0.04-0.02x 0.05-0.04=-0.02x+0.01x 0.01=-0.01x x=-1

OpenStudy (anonymous):

my book says the answer is 3 ?

OpenStudy (lalaly):

wait then let me check my calculations

OpenStudy (lalaly):

my answer is right, recheck if u wrote the question right

OpenStudy (anonymous):

the book says that i guess ill just tell the profesor the book has the wrong answer :)